The perimeter of a rectangle is 98 feet. The ratio of its length to its width s 5:2. Find the area. How do you do this problem!?!?

Answer by TimothyLamb(4379)

---
x = length
y = width
---
2x + 2y = 98
x/y = 5/2
---
put the system of linear equations into standard form:
---
2x + 2y = 98
x = (5/2)y
---
2x + 2y = 98
x - 2.5y = 0
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = length = 35 ft
y = width = 14 ft
---
system matrix
2 2
1 -2.5
---
inverse of system matrix
0.35714286 0.28571429
0.14285714 -0.28571429
---
determinant of system matrix = -7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php